OpenStudy (anonymous):

If a school has 30% students with a dog, 20% with a cat, 10% another pet, and 40% no pets, how can a random number table find the experimental probability that a group of 5 students will have at least 2 dogs? I will make you a fan if you help me

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

im really confused with this.. check this topic it should explain it a little better for you. @doglover789 http://openstudy.com/study#/updates/51578ed8e4b07077e0c06f7e

OpenStudy (anonymous):

Thank you so much for the link...it helped

OpenStudy (anonymous):

no problem :) @doglover789

OpenStudy (anonymous):

I will help you answer your questions if you have any :)

hero (hero):

\[\frac{30}{100} = \frac{3}{10} = \frac{1.5}{5} \approx \frac{2}{5}\] Since you can't have 1.5 dogs you have to round up to the nearest integer.

OpenStudy (anonymous):

Thank you!

hero (hero):

yw

OpenStudy (anonymous):